OpenStudy (anonymous):

WILL GIVE MEDAL!!! PLEASE HELP!! M=y2-y1 (over) x2-x1 solve for y2

OpenStudy (mathstudent55):

\(\large m = \dfrac{y_2 - y_1}{x_2 - x_1} \)

OpenStudy (mathstudent55):

We are solving for \(y_2\). Since \(y_2\) is on the right side, let's start by switching sides. \(\large \dfrac{y_2 - y_1}{x_2 - x_1} = m\)

OpenStudy (anonymous):

okay

OpenStudy (mathstudent55):

There is no \(y_2\) in the denominator, so we can multiply both sides by the denominator of the left side.

OpenStudy (mathstudent55):

\(\large \dfrac{y_2 - y_1}{\cancel{x_2 - x_1}} \times (\cancel{x_2 - x_1}) = m\times (x_2 - x_1)\) \(\large y_2 - y_1 = m(x_2 - x_1)\) Now you only need one more step to finish.

OpenStudy (mathstudent55):

Since you want \(y_2\) alone, you need to add \(y_1\) to both sides.

OpenStudy (mathstudent55):

\(\large y_2 \cancel{- y_1} + \cancel{y_1}= m(x_2 - x_1) + y_1\) \(\large y_2 = m(x_2 - x_1) + y_1\)